monkey69:

What is the equation, in slope-intercept form, of the line that passes through (0, 5) and has a slope of −1? (5 points) a y = −x − 5 b y = x + 5 c y = −x + 5 d y = x − 5

Mercury:

start with the generic form y = mx + b you are given the slope (m) as -1. you are also given a point (0,5) so you can plug in m = -1, x = 0, and y = 5, and solve for b. can you try doing that first?

monkey69:

i cant

Mercury:

well, lets try plugging them in: 5 = (-1)(0) + b what does b equal?

monkey69:

i need the answer pls

monkey69:

im new to this

Mercury:

Please keep in mind we are an educational site and we strive to each, rather than to give out the answer. 5 = (-1)(0) + b can you try multiplying (-1) * 0?

monkey69:

0

Mercury:

good! so 5 = (-1)(0) + b becomes 5 = 0 + b what does b equal now?

monkey69:

5?

Mercury:

awesome! so going back to: y = mx + b we just fill in our m and b values (m = -1 and b = 5) to get y = (-1)x + 5, which is also equal to y = -x + 5 (answer choice c)

monkey69:

thank you